FIFO and LIFO Costs Under Perpetual Inventory System

The following units of an item were available for sale during the year:

Beginning inventory 31 units at $41
Sale 12 units at $59
First purchase 25 units at $42
Sale 7 units at $60
Second purchase 28 units at $45
Sale 30 units at $62

The firm uses the perpetual inventory system, and there are 35 units of the item on hand at the end of the year.

a. What is the total cost of the ending inventory according to FIFO?
$

b. What is the total cost of the ending inventory according to LIFO?
$
Answer #1
FIFO
Date Receipts Issue ( Cost of goods sold) Balance
qty price total qty price total qty price total
1 0 0 31 41 1271
2 0 12 41 492 19 41 779
3 25 42 1050 7 41 287 12 41 492
0 0 25 42 1050
4 28 45 1260 12 41 492 7 42 294
0 18 42 756 28 45 1260
2310 2027
Endig inventory = 294+1260 = 1554

ending inventory = 294+1260 = 1554

LIFO
Date Receipts Issue ( Cost of goods sold) Balance
qty price total qty price total qty price total
1 0 0 31 41 1271
2 0 12 41 492 19 41 779
3 25 42 1050 7 42 294 19 41 779
0 0 18 42 756
4 28 45 1260 2 42 84 19 41 779
0 28 45 1260 16 42 672
2310 2130
ENDING INVENTORY = 779+672 = 1451